2. EXPECTED INTEREST RATE The real risk-free rate is 3 %. Inflation is expected to be 2 % this year and 4 % during the next 2 years. Assume that the maturity risk premium is zero. What is the yield on 2-year Treasury securities? What is the yield on 3 -year Treasury securities?


3. MATURITY RISK PREMIUM The real risk-free rate is 3 %, and inflation is expected to be 3 % for the next 2 years. A 2-year Treasury security yields 6.2 %. What is the maturity risk premium for the 2-year security?


4. PRESENT VALUE OF A CASH FLOW STREAM An investment will pay $ 100 at the end of each of the next 3 years, $ 200 at the end of Year 4 and $ 300 at the end of Year 5 . If other investments of equal risk earn 8 % annually, what is its present value?


5. PRESENT VALUE OF A PERPETUITY What is the present value of a $ 100 perpetuity if the interest rate is 7 % ? If interest rates doubled to 14 %, what would its present value be?


6. EXPECTATIONS THEORY One-year Treasury securities yield 5 %. The market anticipates that 1 year from now, 1-year Treasury securities will yield 6 %. If the pure expectations theory is correct, what is the yield today for 2-year Treasury securities?
Answer #1

2.
a)
=3%+(2%+4%)/2=6%

b)
=3%+(2%+4%*2)/3=6.33%

3.
=(6.2%-3%-3%)=0.2%

4.
=100/1.08+100/1.08^2+100/1.08^3+200/1.08^4+300/1.08^5=608.8906284

5.
a)
=100/7%=1428.57142857143

b)
=100/14%=714.285714285714

6.
=(1.05*1.06)^(1/2)-1=5.50%